# Work performed during an internship at Baidu

### v1

@@ Abstract
Medical image segmentation models adapting to new tasks in a training-free manner through in-context learning is an exciting advancement. Universal segmentation models aim to generalize across the diverse modality of medical images, yet their effectiveness often diminishes when applied to out-of-distribution (OOD) data modalities and tasks, requiring intricate fine-tuning of model for optimal performance. For addressing this challenge, we introduce SegICL, a novel approach leveraging In-Context Learning (ICL) for image segmentation. Unlike existing methods, SegICL has the capability to employ text-guided segmentation and conduct in-context learning with a small set of image-mask pairs, eliminating the need for training the model from scratch or fine-tuning for OOD tasks (including OOD modality and dataset). Extensive experimental validation of SegICL demonstrates a positive correlation between the number of prompt samples and segmentation performance on OOD modalities and tasks. This indicates that SegICL effectively address new segmentation tasks based on contextual information. Additionally, SegICL also exhibits comparable segmentation performance to mainstream models on OOD and in-distribution tasks. Our code will be released after paper review.
Keywords:
Medical Image Segmentation Large Multi-Modal Model Train-free Zero/Few Shot Learning
@@ 1 Introduction
Medical image segmentation is a crucial component of biomedical image analysis and has become a vital tool in medical diagnosis and health monitoring. Medical image segmentation models have evolved from initial convolutional neural network (CNN) structures[30, 48, 8] to the current transformer architectures[35, 15, 6, 42]. Preliminary research outcomes have also been achieved in areas such as efficient data utilization[24, 3] and the design of universal models[5, 22, 26]. With the advancements in this area, the efficient utilization of models for high-quality, multi-modal, and learnable medical image segmentation has emerged as a research hotspot.
@@
Figure 1: The example of inference pipeline of SegICL. The upper part demonstrates SegICL can following text instructions. The lower part demonstrates SegICL can segment OOD data with a few image-mask pairs and text instructions.
In medical image segmentation, two key research directions emerge: handling OOD tasks, and segmentation of regions of interest. Existing methods for the above two issues can be categorized into two types: specialized models designed for specific tasks and universal segmentation methods.
Firstly, specialized model for OOD tasks, like few-shot learning[39, 17, 34] paradigm aimed at improving the generalization of the model to new data with a small number of examples. Another approach involves enhancing the model’s generalization performance through data augmentation and pseudo-labels[3, 4, 23, 45, 36, 24, 43]. There are also universal segmentation models, designed to handle segmentation tasks across multiple medical modalities. However, specialized model’s generalization performance is mediocre, particularly when facing cross-modal or new tasks. While universal segmentation models exhibit good generalization performance, they are unable to handle OOD tasks. However, given the rapid progress in the medical field, if a completely new modality of medical imaging is utilized, all the above models would require retraining or fine-tuning.
Secondly, for segmentation of regions of interest, one of the existing solutions is to build a universal segmentation model based on SAM[19]. It utilizes geometric information such as points and boxes as prompts to guide the model in segmentation. Consider an interesting question: How can an annotator lacking professional medical training quickly annotate organs they are unfamiliar with on medical images they don’t understand? For instance, given a CT imaging, how can an annotator who lacks knowledge about the shape and approximate location of the ’spleen’ quickly annotate its position?
As illustrated in Fig. 1, SegICL tackles these two issue through a text-guided interactive in-context learning paradigm. Unlike existing specialized model and universal segmentation model that demand retrain or fine-tuning to improve the segmentation performance on OOD tasks, SegICL is a train-free method which segments regions of interest by integrating text instructions and a few image-mask pairs. By leveraging a few number of image-mask pairs, SegICL achieves remarkable performance improvement on OOD tasks.
In summary, our contributions are as follows:
  • For the first time, we propose a universal in-context learning paradigm for medical image segmentation. In this learning paradigm, the model can learn a in-context visual representation and follow text instrctions.
  • SegICL utilizes in-context learning to solve segmentation of OOD tasks issues, tackles region-of-interest segmentation through text instructions.
@@
  • Extensive experiments have demonstrated that SegICL exhibits impressive performance across multiple medical image segmentation datasets, showcasing competitive capabilities comparable to other SOTA methods.
@@ 2.1 Universal Medical Image Segmentation
With the rise of universal vision models such as CLIP[28], SAM[20] and SegGPT[38], the field of biomedical image segmentation has also witnessed a surge in research on universal medical segmentation models.
These studies can be classified into two types. One type involves fine-tuning or replacing new modules based on existing general vision models (such as SAM)[46, 40, 7, 9, 22]. For example, SAMed[46] adopts a low-rank-based fine-tuning strategy (LoRA) on top of SAM, fine-tuning the SAM image encoder on labeled medical image segmentation datasets. Another example is Med-SA[40], which differs from fine-tuning the SAM model by proposing the Medical SAM Adapter (Med-SA), incorporating domain-specific medical knowledge into the segmentation model. The other type involves training specialized universal medical segmentation models from scratch based on large-scale datasets. For instance, UniverSeg[5] utilizes a novel CrossBlock mechanism to generate accurate segmentation maps without requiring additional training, achieving generalization to new tasks.
Our approach differs from the methods mentioned above. The core of our proposed paradigm lies in a multi-modal input-driven segmentation method that incorporates in-context learning based on textual instructions. None of the aforementioned methods can simultaneously achieve both text-guided segmentation and context-aware segmentation, which are the two core aspects of our approach.
@@ 2.2 Learning from limited data samples
Few-shot medical image segmentation[39, 17, 34, 33, 2, 37] is designed to efficiently address the scarcity of learning and generalizing from a limited number of samples. In the context of segmentation tasks, few-shot learning is introduced to segment unseen classes with the support of a limited amount of labeled data.
VQNet[45] introduces a VQ learning mechanism with Grid-Format VQ (GFVQ), Self-Organizing VQ (SOVQ), and Residual-Guided VQ (ROVQ) to medical MR imaging segmentation. SE-Net[31] incorporate few-shot learning for segmenting abdominal organs in CT images. RPNet[34] proposed a method that involves a context relation encoder (CRE) for capturing local relation features and a recurrent mask refinement module. This module iteratively refines the segmentation mask by utilizing the CRE and a prototypical network.
SegICL’s motivation aligns with few-shot learning, aiming to learn OOD tasks or adapt to OOD datasets or modality with minimal annotated data. However, unlike the aforementioned methods, these approaches exhibit poorer generalization performance, making them challenging to apply across modalities and tasks. Moreover, when confronted with OOD tasks from different modalities, they struggle to segment without the need for retraining. In contrast, the proposed paradigm possesses the capability of multi-modal context learning, enabling it to achieve learning without retraining when facing OOD tasks.
@@ 3 Methodology
3.1 Learning Paradigm of SegICL
The learning paradigm of SegICL aims to utilize information from multi-modal input, enabling the model to model and extract implicit data correlations in multi-modal data. Additionally, it empowers the model with the ability to perform segmentation using text-guided methods.
@@ 3.1 Learning Paradigm of SegICL
The learning paradigm of SegICL (See Fig. 2) can be represented as Eqs. 1, 2 and 3: where $X_{img}$ and $X_{text}$ represents the input including images and texts respectively, $Enc$ is a large vision language model. After passing the multi-modal input through $Enc$, it obtains a hidden variable ($H$). Then, using a projector to align the $H$ of the encoder with the feature space of the decoder, obtaining the encoded result state ($S$). Finally, the state is fed into the $Dec_{img}$ for decoding, resulting in the final output mask.
$M^{\prime}=Dec_{img}(S_{img})$ (3)
Figure 3: Training pipeline of SegICL. Module A is a image decoder ($Dec_{img}$), and Module B is a Multi-modal encoder ($Enc$). Interleaved multi-modal text and image input undergoes encoding in Module B (get the hidden variable $H_{img}$), and the projected features ($S_{img}$) is then passed to Module A for decoding, ultimately generating the corresponding mask (${}^{\prime}$).
@@
The $Enc$ is a trainable function that accepts multi-modal input and regresses to a hidden state features. The $Dec$ can be any image generation model, decoding the final prediction from above state. The $Enc$’s supervision signal comes from the regression loss between the encoded result state and the condition encoder (See Fig. 3). The $Dec_{img}$’s supervision signal comes from the regression loss between the generated image and the ground truth masks.
3.2 SegICL
Building upon the aforementioned learning paradigm, to our best knowledge, SegICL is the first text-guidence in-context learning framework for medical image segmentation.
@@ 3.2 SegICL
This framework (See Fig. 3) use a LLM implementing the $Enc$ to model the input multi-modal data. Additionally, it includes a diffusion model ($Dec_{img}$) to decode the state vector outputted by the encoder, yielding the final prediction. A 4-layer Multi-Layer Perceptron (MLP) serves as the shared condition encoder to provide regression targets for the encoder model.
$M^{\prime}=Dec_{img}(Proj((Enc(X_{img},X_{text}))))$ (4)
@@
The forward computation of the framework is represented by the equation above (see Eq. 4). This framework initially takes arbitrary interleaved textual and visual data as input. After processing through a visual encoder and a text encoder, it generates interleaved tokens of text and image (See Fig. 3 module B). These tokens serve as input for the multi-modal model, generating latent variables within a transformer structure rich in causal self-attention mechanisms.
Since the encoder and decoder are decoupled, a projector is needed to map the model’s feature space. After the tokens pass through the projector, a state vector is generated and made available for the decoder. The decoder model utilizes a diffusion model to ensure the quality of the generated images. When the input state vector is provided to the decoder model, it introduces random noise and uses the state vector as a condition to predict the output mask (See Fig. 3 module A).
SegICL’s diffusion decoder is trained using a substantial amount of ground truth mask images. Its input includes random noise and a conditional vector (obtained by encoding mask images through the condition encoder). The output is a mask image, with the goal of reconstructing the original image using the vector provided by the shared condition encoder.
@@ 3.3 Multi-modal Encoder
In contrast to existing multi-modal large models[33, 12, 2, 1], the multi-modal encoder used in this paper aims to regress the supervision signals provided by the condition encoder. Therefore, in the model design, a lightweight condition encoder is incorporated for supervised fine-tuning of the multi-modal large model.
Multi-modal encoder’s network architecture of the encoder is depicted in Fig. 2 module A. The visual encoder utilize CLIP ViT-big-G[28] as the base model and added a single-layer randomly initialized cross-attention layer on it. Visual encoder trained with 1.5 billion text-image data, the model supports an input image resolution of 448$\times$448, significantly enhancing its performance for fine-grained segmentation tasks. The language base model uses the pre-trained Qwen-7B model, contributing to a substantial reduction in the difficulty of transfer training, thanks to this excellent open-source work.
$\mathcal{L}=\frac{1}{n}\sum_{i=1}^{n}\left(\text{S}_{i}-\hat{\text{C}}_{i}\right)^{2}$ (5)
The condition encoder implemented with a same 4-layer MLP structure as the diffusion’s condition encoder, ensures a lightweight network architecture to maintain model performance while significantly simplifying training complexity and parameter volume. We have standardized the input length for images, encoding each image into 256 tokens. Consequently, the language model’s final output is a token sequence with this length. Inspired by LLaVA, we employed a simple projection layer to transform the feature space from the language model to the same feature space as the condition encoder. Ground truth masks have already undergone encoding by the condition encoder, producing a conditional variable with a length of 256. By infinitely narrowing the gap between the output of the teacher network and the output of the language model, we can ensure the final generation quality of the decoder. Therefore, the loss function for the multi-modal encoder is as shown above Eq. 5, using the MSE loss function to efficiently model the latent mapping between result state $S_{i}$ and $\hat{\text{C}}_{i}$, where $S_{i}$ represent projected image token, $\hat{\text{C}}_{i}$ represent the ground truth masks encoded by condition encoder.
In addition, the exploration involves utilizing CLIP-ViT-big-G as the condition encoder and fitting multi-modal encoder’s output with its output. Experimental comparisons will be presented in the appendix to elucidate our choice.
@@ 3.4 Image Decoder
The image decoder can be any model capable of performing image generation tasks. In this paper, ControlNet[47] is chosen as the image decoder, and its model structure can be referred to in Fig. 3, Module B. In this context, the task of the image decoder is to accept a set of conditional vectors and generate images based on these vectors.
$\mathcal{L}=\mathbb{E}_{m_{0},t,c_{k},c_{\text{t}},\epsilon\sim\mathcal{N}(0,1)}\Big{[}\|\epsilon-\epsilon_{\theta}(m_{t},t,c_{k},c_{\text{t}}))\|_{2}^{2}\Big{]}$ (6)
We utilized the pre-training parameters of the Stable Diffusion 1.5B[29] model to ensure the image decoder could be trained with relatively less data while achieving excellent image generation results. The training process for the image decoder can be conceptualized as the recovery of input image features from the conditioned vectors. Initially, the input image is fed into the designated conditional encoder, extracting features from the input image to generate conditional variables. Subsequently, the conditional variables are injected into the backbone network using cross-attention. The loss is calculated by comparing the noise to optimize the entire image encoder during the training process (See Eq. 6). A mask image as input (${m}_{0}$), then diffusion algorithms progressively introduce noise to the image, generating a noisy image (${m}_{t}$). Subsequently, with the keyword prompt (${c}_{k}$) and the feature vector output from the conditional model(${c}_{t}$) as auxiliary, a function ($\epsilon_{\theta}$) is employed to predict noise.

### v4

# SegICL: A Multimodal In-context Learning Framework for Enhanced Segmentation in Medical Imaging

@@ Abstract
In the field of medical image segmentation, tackling Out-of-Distribution (OOD) segmentation tasks in a cost-effective manner remains a significant challenge. Universal segmentation models is a solution, which aim to generalize across the diverse modality of medical images, yet their effectiveness often diminishes when applied to OOD data modalities and tasks, requiring intricate fine-tuning of model for optimal performance. Few-shot learning segmentation methods are typically designed for specific modalities of data and cannot be directly transferred for use with another modality. Therefore, we introduce SegICL, a novel approach leveraging In-Context Learning (ICL) for image segmentation. Unlike existing methods, SegICL has the capability to employ text-guided segmentation and conduct in-context learning with a small set of image-mask pairs, eliminating the need for training the model from scratch or fine-tuning for OOD tasks (including OOD modality and dataset). Extensive experimental demonstrates a positive correlation between the number of shots and segmentation performance on OOD tasks. The performance of segmentation when provided thre-shots is approximately 1.5 times better than the performance in a zero-shot setting. This indicates that SegICL effectively address new segmentation tasks based on contextual information. Additionally, SegICL also exhibits comparable performance to mainstream models on OOD and in-distribution tasks. Our code will be released after paper review.
@@ 1 Introduction
Medical imaging segmentation is a crucial component of biomedical image analysis and has become a vital tool in medical diagnosis and health monitoring. It have evolved from initial CNN structuresa42 ; zhou2018unet++ ; a3 to the vision transformera4 ; a5 ; a6 ; a20 . Preliminary research outcomes have also been achieved in areas such as efficient data utilizationa7 ; a8 and the design of universal modelsa9 ; a10 ; a12 . With the advancements in this area, researchers seek a framework with strong generalization performance, low training and maintenance costs, and user-friendly.
In the field of medical imaging segmentation, enhancing model generalization performance on OOD tasks and reducing develop cost remains a challenge. Traditional methods such as Few-Shot Learning (FSL) aim to improve model generalization with a small number of shots a16 ; a17 ; a18 . However, specific FSL models often excel only within specific modalities, and tend to underperform when applied to other types of medical imaging datasets. Semi-supervised learning approaches use data augmentation or pseudo-labeling techniques a8 ; a11 ; a13 ; a14 ; a19 ; a7 ; a15 to reduce the costs associated with data annotation. However, these methods usually require fine-tuning or retraining when facing to new tasks, which incurs additional costs and effort. ICL methods offers a promising alternative, as it can effectively handle Out-of-Distribution (OOD) tasks in a cost-efficient manner. It do not require retraining to adapt to new tasks, which significantly lowers maintenance costs. Despite their potential, existing ICL methods have primarily been explored in domains such as language tasks a63 ; a64 ; a65 ; a60 ; a61 , image generation a66 ; a67 , image segmentation and grounding a50 ; a70 ; a71 , multimodal understanding a72 ; a23 ; a74 ; a76 ; a62 , and embodied vision a75 . As far as we know, there have been limited attempts to apply ICL to fine-grained medical perception tasks like segmentation. From the perspective of user-friendly interaction, one notable method involves building a universal segmentation model based on the SAMa28 . This model uses geometric information such as points and boxes as prompts to guide segmentation. However, this approach can be less user-friendly if users lack the medical knowledge to accurately identify geometric locations of certain organs.
In summary, while FSL methods can somewhat alleviate OOD task issues, they struggle with cross-modal datasets. Semi-supervised methods help reduce development costs by minimizing data annotation requirements but still need fine-tuning or retraining for new tasks. Existing ICL methods show strong generalization and adaptability without retraining, yet they have not been adequately explored for fine-grained medical tasks.
@@
Figure 1: The example of inference pipeline of SegICL. The left part demonstrates SegICL can following text instructions. The right part demonstrates SegICL can segment OOD data with a few image-mask pairs and text instructions.
To address these limitations, we propose SegICL, the first multimodal interactive in-context learning paradigm for medical segmentation, as shown in Fig. 1. SegICL addresses OOD tasks by performing ICL with few shots, making it suitable for segmenting various image modalities and avoiding the poor transfer performance of FSL. In terms of cost reduction and adaption of new tasks, SegICL adapts to new tasks with a training-free manner, thereby lowering maintenance and training costs associated with model fine-tuning or retraining. For user-friendliness, SegICL leverages text to guide the segmentation of target areas, which reduces the adaptation threshold for users.
@@
Our contributions are as follows. Firstly, we propose a ICL paradigm for medical image segmentation, which can learn a in-context visual representation and follow multimodal instructions. Secondly, SegICL utilizes ICL to solve segmentation of OOD tasks, tackles region-of-interest segmentation through user-friendly instructions. Thirdly, extensive experiments have demonstrated that SegICL exhibits impressive performance across multiple medical image segmentation datasets, showcasing competitive capabilities comparable to other SOTA methods.
@@ 2.1 Universal medical image Segmentation
With the rise of universal vision models such as CLIPa29 , SAMa45 and SegGPTa50 , the field of biomedical image segmentation has also witnessed a surge in research on this topic. These studies can be classified into two types. One method involves fine-tuning or replacing new modules based on existing general vision models (such as SAM)a46 ; a47 ; a48 ; a49 ; a10 . For example, SAMeda46 adopts a low-rank-based fine-tuning strategy (LoRA) on top of SAM, fine-tuning the SAM image encoder on labeled medical image segmentation datasets. Another example is Med-SAa47 , which differs from fine-tuning the SAM model by proposing the Medical SAM Adapter (Med-SA), incorporating domain-specific medical knowledge into the segmentation model. The other type involves training specialized universal medical segmentation models from scratch based on large-scale datasets. For instance, UniverSega9 utilizes a novel CrossBlock mechanism to generate accurate segmentation mask without requiring additional training, achieving generalization to new tasks.
Our approach differs from the methods mentioned above. The core of our proposed paradigm lies in a multimodal input-driven segmentation method that incorporates in-context learning based on textual instructions. None of the aforementioned methods can simultaneously achieve both text-guided segmentation and context-aware segmentation, which are the two core aspects of our approach.
@@ 2.2 Learning from limited data samples
Few-shot medical image segmentationa16 ; a17 ; a18 ; a22 ; a24 ; a38 is designed to efficiently address the scarcity of learning and generalizing from a limited number of shots. In the context of segmentation tasks, few-shot learning is introduced to segment unseen classes with the limited dataset. VQNeta14 introduces a VQ learning mechanism with Grid-Format VQ (GFVQ), Self-Organizing VQ (SOVQ), and Residual-Guided VQ (ROVQ) to medical MRI image segmentation. SE-Neta39 incorporate few-shot learning for segmenting abdominal organs in CT images.
SegICL’s motivation aligns with few-shot learning, aiming to learn OOD tasks or adapt to OOD datasets with minimal annotated data. However, these approaches exhibit poorer transfer performance, making them challenging to apply across modalities and tasks. Moreover, when confronted with OOD tasks from different modalities, they struggle to complete the segmentation accurately. In contrast, the proposed paradigm possesses the capability of multimodal ICL, enabling it to wrok without retraining when facing OOD tasks.
@@ 3 Methodology
3.1 In-Context Learning Paradigm of SegICL
The learning paradigm of SegICL aims to utilize information from multimodal input, enabling to extract implicit data correlations in multimodal data. Additionally, it empowers the model with the ability to perform segmentation using text-guided methods.
@@ 3.1 In-Context Learning Paradigm of SegICL
The learning paradigm of SegICL (See Fig.2) can be represented as eq.1-eq.3: where $X_{img}$ and $X_{text}$ represents the input including images and texts respectively, $Enc$ is a large language model (LLM). After passing the multimodal input through $Enc$, it obtains a hidden variable ($H$). Then, using a projector to align the $H$ of the encoder with the feature space of the decoder, obtaining the encoded result state ($S$). Finally, the state is fed into the $Img_{Dec}$ for decoding, resulting in the final output mask.
$M^{\prime}=Img_{Dec}(S)$ (3)
@@
The $Enc$ is a trainable function that accepts multimodal input and regresses to a hidden state vector. The $Img_{Dec}$ can be any image generation model, decoding the final prediction from above state. The $Enc$’s supervision signal comes from the regression loss between the encoded result state and the condition encoder (See Fig.3). The $Img_{Dec}$’s supervision signal comes from the regression loss between the generated image and the ground truth masks.
Figure 3: Training pipeline of SegICL. Module A is a image decoder, and Module B is a multimodal encoder. Interleaved multimodal text and image input undergoes encoding in Module B (get the hidden variable H ), and the projected data (get the result state S) is then passed to Module A for decoding, ultimately generating the corresponding mask.
3.2 Architecture
@@
Building upon the aforementioned learning paradigm, we propose the first universal text-guidence in-context learning medical segmentation framework SegICL. This framework (See Fig.3) use a LLM implementing the $Enc$ to model the input multimodal data. Additionally, it includes a diffusion model ($Img_{Dec}$) to decode the state vector outputted by the encoder, yielding the final prediction. A MLP serves as the shared condition encoder to provide regression targets for the encoder model.
$M^{\prime}=Img_{Dec}(Proj((Enc(X_{img},X_{text}))))$ (4)
@@
The forward computation of the framework is represented by the equation above (see Fig.4). This framework initially takes arbitrary interleaved textual and visual data as input. After processing through a visual encoder and a text encoder, it generates interleaved tokens of text and image (See Fig.3 module B). These tokens serve as input for the multimodal model, generating latent variables within a transformer structure rich in causal self-attention mechanisms. Since the encoder and decoder are decoupled, a projector is needed to map the model’s feature space. After the tokens pass through the projector, a state vector is generated and made available for the decoder. The decoder model utilizes a diffusion model to ensure the quality of the generated images. When the input state vector is provided to the decoder model, it introduces random noise and uses the state vector as a condition to predict the output mask (See Fig.3 module A). SegICL’s $Img_{Dec}$ is trained using ground truth mask images. Its input includes random noise and a conditional vector (obtained by encoding mask images through the condition encoder). The output is a segmentation mask, with the goal of reconstructing the original image using the vector provided by the shared condition encoder.
@@ 3.3 Multi-modal Encoder
In contrast to existing multimodal large modelsa22 ; a23 ; a24 ; a25 , the multimodal encoder used in this paper aims to regress the supervision signals provided by the condition encoder. Therefore, a lightweight condition encoder is incorporated for supervised fine-tuning of the multimodal Encoder.
$\mathcal{L}=\frac{1}{n}\sum_{i=1}^{n}\left(\text{H}_{i}-\hat{\text{C}}_{i}\right)^{2}$ (5)
The input $img\in\mathbb{R}^{H\times W\times 3}$ and $text\in\mathbb{R}^{\ Len\times 1}$ is processed by $Enc$ to obtain $M\in\mathbb{R}^{C_{hs}\times N_{v}}$, where H represents a multimodal input token vector ,$C_{hs}$ is the hidden size in the multimodal Encoder. M is then fed into the multimodal encoder and get the image feature vectors $H\in\mathbb{R}^{C_{hs}\times N_{v}}$ and text feature vectors $\hat{text}\in\mathbb{R}^{C_{hs}\times N_{v}}$. Finally, the $\hat{text}$ are sent to the language head for decoding, while the $H$ are sent to the image decoder for decoding (See Fig.3).
The language base model uses the pre-trained Qwen-7B model. The condition encoder implemented with a same MLP structure as the diffusion’s condition encoder. By infinitely narrowing the gap between the output of the condition encoder and the output of the multimodal encoder, we can ensure the final generation quality of the decoder. Therefore, the loss function use the MSE loss(See eq.5) function to efficiently model the latent mapping between image feature $H$ and condition $C$.
@@ 3.4 Image Decoder
The $Img_{Dec}$ can be any model capable of performing image generation tasks. In this paper, ControlNeta31 is chosen as the $Img_{Dec}$, and its model structure can be referred to in Fig.3, Module B. In this context, the task of the image decoder is to accept a set of conditional vectors and generate images based on these vectors.
$\mathcal{L}=\mathbb{E}_{\bm{m}_{0},\bm{t},\bm{c}_{k},\bm{c}_{\text{t}},\epsilon\sim\mathcal{N}(0,1)}\Big{[}\|\epsilon-\epsilon_{\theta}(\bm{m}_{t},\bm{t},\bm{c}_{k},\bm{c}_{\text{t}}))\|_{2}^{2}\Big{]}$ (6)
A segmentation mask ${m}_{0}\in\mathbb{R}^{H\times W\times 3}$ as input, then pass VAE encoder and diffusion algorithms progressively introduce noise to the image, generating a latent space image ${m}_{t}\in\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times 4}$. Subsequently, with the keyword prompt ${c}_{k}\in\mathbb{R}^{1\times 77\times dim_{p}}$ and the feature vector output from the conditional model ${c}_{t}\in\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times dim_{c}}$ as auxiliary, a function $\epsilon_{\theta}$ is employed to predict noise. $dim_{p}$ and $dim_{c}$ is the the dimensions of the features. We utilized the pre-training parameters of the SD 1.5Ba30 to ensure the image decoder could be trained with relatively less data while achieving excellent image generation results. The training process can be conceptualized as the recovery of input image features from the conditioned vectors. Initially, the input image is fed into the designated conditional encoder, extracting features from the input image to generate conditional variables. Subsequently, the conditional variables are injected into the backbone network using cross-attention. The loss is calculated by comparing the noise to optimize the entire image encoder during the training process.